OpenStudy (anonymous):

write an equation of the line containing a given point and parallel to the given line. Express you answer in the given form y=mx+b (7,8);x+3y=2

OpenStudy (anonymous):

x+3y=2=>y=(-1/3)*x+(2/3) y=(-1/3)*x+b 8=(-1/3)*7+b b=8+(7/3)=31/3 y=(-1/3)*x+(31/3)